# Robust Plan Evaluation based on Approximate Probabilistic Machine Learning

### v2

@@ 3.2.2. Risk Quantification
Suboptimality Risk As stated in the problem formulation with Approach 1, the goal of robust query optimization is to find a plan with minimum risk of suboptimality. This can be formalized as the likelihood of the plan being suboptimal compared with the alternative plan(s). To explain this, let us revisit Example 3.2. We assume that the cost of the optimal plans in all scenarios is 8, and it is 10 for the alternative plans. The difference between the scenarios is the standard deviation of the probability distributions. Let us say the cost for the optimal plan (plan X) is denoted by $C(X)\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and the cost for the alternative plan (plan Y) is denoted by $C(Y)\sim\mathcal{N}(\mu_{y},\sigma_{y}^{2})$. Also, let us assume the covariance between $C(X)$ and $C(Y)$ is denoted as $\text{Cov}(x,y)$. Then the probability of the alternative plan being cheaper than the optimal plan is denoted by:
@@
Computation Overheads In order to minimize the compilation overheads, the computation of SOR can be vectorized, such that at each step of comparing several plan fragments $R(p_{i},p_{j})$ is computed as a matrix $R_{n\times n}$ where $n$ is the number of plans to be compared and $R_{i,j}=R(p_{i},p_{j})$. To this end, we define $D_{n\times n}$ where $D_{i,j}=\mu_{i}-\mu_{j}$ and $S_{n\times n}$ where $S_{i,j}=\sqrt{\sigma_{i}^{2}+\sigma_{j}^{2}}$. Then the z-scores for various comparing different pairs of plans can be computed as:
(14) $Z_{n\times n}=-{D_{n\times n}}\oslash{S_{n\times n}}\quad$
@@
An accurate estimation of the risk associated with selecting the plan $X$ over the plan $Y$ would require the probability distribution of the cost for each plan, as well as the pairwise joint distribution $C(X)\cdot C(Y)$. Assuming the independence between $C(X)$ and $C(Y)$ results in overestimating the risk by increasing the variance of $C(X)-C(Y)$. We assume independence for two reasons: a) the computation overheads would be significant and b) even with the independence assumption made, the risk quantification provides significant improvements. The effectiveness of the simplified quantification, as demonstrated in the experimental study, can be explained by the fact that the overestimation effect for all pairs of plans is consistent, effectively canceling each other out.
The Probability Distribution The cost or execution time are not inherently normally distributed variables. However, they can be transformed such that they approximately conform to a normal distribution. This is handled in the preprocessing phase as explained in the experimental study section.
3.3. Risk-aware Plan Evaluation
The classic optimizers select a plan from a set of plans solely based on the expected cost. By doing so, they assume that the estimated expected costs have zero uncertainties. In contrast, when the uncertainties, and, thus, risk, are quantified, a new family of plan evaluation and selection methods can be devised that account for them. We propose three risk-aware plan evaluation strategies. While the first two strategies use risk measures directly for selecting plans, the third uses risk to prune the search space. We design algorithms that implement these strategies.
3.3.1. Plan Selection by Suboptimality Risk
Plan selection by suboptimality risk follows the first approach to the problem formulation where the problem is defined a risk minimization one. As such, the suboptimality risk is computed for every plan in the search space, and the one with the minimum risk is selected, based on the risk quantification method suggested in section 3.2. Given a set $\mathcal{P}=\{p_{i}|\forall i\in\{1,\ldots,S\}\}$ of $S$ plans, it computes the suboptimality risk $\text{SOR}(p_{i})$ for $p_{i}$. The plan with minimum $\text{SOR}(p_{i})$ is selected. The algorithm can use either plan or estimation risk, or a combination of the two.
@@ 3.3.1. Plan Selection by Suboptimality Risk
3:  Initialize $P^{*}\leftarrow\emptyset$
4:  Initialize $SOR\leftarrow\emptyset$ an empty set for subopt risks
5:  for $p_{i}\in\mathcal{P}$ do
6:     for $p_{j}\in\mathcal{P}$ do
7:        if $j\neq i$ then
8:           $C_{p_{i}}-C_{p_{j}}\sim N(\mu_{i}-\mu_{j},\sigma_{i}^{2}+\sigma_{j}^{2}-2\text{Cov}(C_{p_{i}},C_{p_{j}}))$
9:           $R(p_{i},p_{j})\colon=P(C_{p_{i}}-C_{p_{j}}>0)$
10:        end if
11:     end for
12:     $R(p_{i})\colon=\frac{1}{S-1}\sum_{j=1,j\neq i}^{S}R(p_{i},p_{j})$
13:     $SOR\leftarrow SOR\cup\{R(p_{i})\}$
14:  end for
15:  $y^{*}\colon=\arg\min_{i\in\{1,\ldots,S\}}SOR$
16:  $P^{*}\leftarrow p_{y^{*}}$
17:  return $P^{*}$
3.3.2. Conservative Plan Selection
The conservative plan selection strategy assumes that the plans have an execution cost that is higher than their estimated cost, and this difference is proportional to the standard deviation of their estimated cost. It eliminates risky plans (that is, plans with high standard deviation) that happen to be low cost. Given a set $\mathcal{P}=\{p_{i}|\forall i\in\{1,\ldots,S\}\}$ of $S$ plans, where each $p_{i}$ has a cost distribution $C(p_{i})\sim\mathcal{N}(\mu_{p_{i}},\sigma_{p_{i}}^{2})$, rather than picking the plan with minimum $\mu_{p_{i}}$, it picks the plan with minimum $\mu_{p_{i}}+f_{s}\sigma_{p_{i}}$, where $f_{s}$ is a parameter tuned on a validation set. The algorithm can be specialized to use either the plan or the estimate risk, or both. This algorithm follows the second approach to the problem formulation where a linear combination of the two objectives $\alpha.\mu_{p_{i}}+(1-\alpha).\sigma_{p_{i}}$ is minimized, where $f_{s}=\frac{1-\alpha}{\alpha}$.
@@ 3.3.2. Conservative Plan Selection
3:  Initialize $P^{*}\leftarrow\emptyset$
4:  Initialize $CC\leftarrow\emptyset$ an empty set for conservative cost estimates
5:  for $p_{i}\in\mathcal{P}$ do
6:     $CC_{i}\colon=\mu_{p_{i}}+f_{s}\cdot\sigma_{p_{i}}$
7:     $CC\leftarrow CC\cup\{CC_{i}\}$
8:  end for
9:  $y^{*}\colon=\arg\min_{i\in\{1,\ldots,S\}}CC$
10:  $P^{*}\leftarrow p_{y^{*}}$
11:  return $P^{*}$
@@ 3.3.3. Search Space Pruning by Plan and Estimation Risks
This strategy prunes the search space to eliminate highly risky plans, based on either the plan or the estimation risks. Algorithm 3 is designed to implement this strategy: given a set $\mathcal{P}=\{p_{i}|\forall i\in\{1,\ldots,S\}\}$ of $S$ plans along with their plan and estimation risks ($R_{p}$ and $R_{e}$ respectively), and fractions of plans to be pruned with the highest plan and estimation risks ($f_{pr}$ and $f_{er}$ respectively), thresholds $\phi_{pr}$ and $\phi_{er}$ are computed for plan and estimation risks, respectively. Then, any plan with either plan or estimation risks above the designated thresholds is eliminated. Parameters $f_{pr}$ and $f_{er}$ are tuned using the samples in the validation set. Choosing the plan from the remaining ones relies on a plan selection strategy.
@@
1:  Inputs: $\mathcal{P}=\{p_{i}\mid\forall i\in\{1,\dots,S\}\}$: a set of plans to be evaluated $R_{e}=\{R_{e}(p_{i})\mid\forall i\in\{1,\dots,S\}\}$: a set of estimation risks $R_{p}=\{R_{p}(p_{i})\mid\forall i\in\{1,\dots,S\}\}$: a set pf plan risks $f_{er}$: a fraction of plans to be pruned by estimation risk $f_{pr}$: a fraction of plans to be pruned by plan risk
2:  Output: $P^{\prime}\subset\mathcal{P}$: the pruned search space
3:  Sort $R_{e}$ to get $R_{es}$
4:  $\text{ind}_{er}=\lfloor|\mathcal{P}|\times f_{er}\rfloor$
5:  $\varphi_{er}=R_{es}[\text{ind}_{er}]$
6:  Sort $R_{p}$ to get $R_{ps}$
7:  $\text{ind}_{\text{pr}}=\lfloor|\mathcal{P}|\times f_{pr}\rfloor$
8:  $\varphi_{pr}=R_{ps}[\text{ind}_{\text{pr}}]$
9:  $P^{\prime}=\mathcal{P}$
10:  for $i=1$ to $S$ do
11:     if $R_{e}(p_{i})>\varphi_{er}$ then
12:        $P^{\prime}=P^{\prime}\setminus\{p_{i}\}$
13:     end if
14:     if $R_{p}(p_{i})>\varphi_{pr}$ then
15:        $P^{\prime}=P^{\prime}\setminus\{p_{i}\}$
16:     end if
17:  end for
18:  return $P^{\prime}$
@@ 4. Risk-aware Learned Cost Model
We propose a novel learned cost model designed to be robust to changes in input samples. It is designed to infer not only the expected execution time, but also the plan, the estimation, and suboptimality risks. Figure 3 illustrates the architecture of the model. The model takes as input queries and plans and predicts the expected execution time along with expected variance. The architecture includes two main modules: The first learns the representation of query and plan combinations; the second takes this representation and learns the query-plan associations and their impact on runtime.
Figure 3. The architecture of the proposed risk-aware learned cost model, including representation learning and estimator components. This architecture allows quantification of model and data uncertainties.
[…]
@@
Query Encoding. For the query representation, we use the query’s join graph to encode information about tables, local predicates, join predicates, aggregations, etc. Similar approaches were used in previous work (Yu et al., 2022). In this approach, we encode table statistics, such as cardinality and selectivity from local predicates (estimated by the optimizer) and correlations between the predicate columns (captured by the optimizer statistics) as node attributes. We also incorporate characteristics of the join predicates, such as the join type, the join predicate operator, the join column skewness and the join selectivity, as edge attributes. The optimizer statistics are extended to capture the join characteristics, too. In addition to the node and edge attributes, we capture high-level characteristics of the join graph, such as the number of tables and joins, and the join graph topology, as global attributes. This representation is agnostic to the plan (join orders and plan operators) that will be used to execute the query.
@@
Plan Encoding. The plan-level information is encoded using a vectorized plan tree that preserves the plan structure (Marcus et al., 2019). Each node in the vectorized tree corresponds to an operator node in the plan tree. Each node captures the operator type, including different access, join, and aggregate operators, using one-hot encoding. In addition, it captures the table(s) included in the subgraph leading to that node using an embedding learned by the Query processor as explained in the following section (see Figure 3).
@@
Representation Learning. The representation learning module takes the query-level and plan-level information as input and produces query embeddings (using a Query Processor) and plan embeddings (using a Plan Processor). While we partially borrow the architecture of the Plan Processor from prior works (Marcus et al., 2019), the architecture of the Query Processor and how that feeds the Plan Processor is novel and supports the robustness of the proposed model, as demonstrated in the experiments.
@@
Query Processor. We use Graph Neural Nets (GNNs) to process the encoded join graph and learn query and table representations. Each GNN layer serves as a message-passing step where each table receives information from its adjacent tables, with each additional layer adding to the depth of the message massing in the graph. This allows learning a representation for each table such that it contains information about its adjacent tables as well as the whole join graph. Additionally, we use mean and max pooling to aggregate the information from all tables and produce a representation for the whole join graph. While various GNN implementations exist, we use TransformerConv (Shi et al., 2021) in our architecture. The TransformerConv uses an attention mechanism (Vaswani et al., 2017) to condition the amount of information that each node receives from its neighboring nodes by the attributes of the target node, the neighboring nodes, and the edges connecting them to the target node. This is a suitable option for processing the join graphs as the amount of data processed through a subgraph should be conditioned by the tables and the joins involved. For example, considering a join graph where a fact table is joined to a few dimension tables; the amount of influence from each dimension table on the final cardinality depends on the characteristics of the tables and the join types and columns.
@@
Plan Processor. The representations of the plans are learned by the Plan Processor, which takes the vectorized plan tree as input and augments each node with the embedding of the corresponding table(s) produced by the Query Processor. If the plan node is an access operator, it is augmented with the embedding of the corresponding table only. If a node is a subgraph involving multiple tables, the embeddings of the corresponding tables are aggregated using mean-pooling before being concatenated to the plan node. The augmented plan tree is then processed through multiple Tree Convolutional Neural Nets (TCNN) layers, ultimately aggregated into a one-dimensional vector using dynamic pooling (Mou et al., 2016). This architecture is proven to be effective in processing query plan trees by prior works (Marcus et al., 2019, 2024; Yu et al., 2022). The embeddings generated by the final layer of this model serves as the representation of the query plan.
@@ 5. Experimental Study
We present an experimental study that evaluates Roq in comparison to the state-of-the-art.
@@ 5.1. Experimental Setup
We use the following benchmarks: The Join Order Benchmark (JOB) (Leis et al., 2015), the Cardinality Estimation Benchmark (CEB) (Negi et al., 2021), and the Decision Support Benchmark (TPC-DS) (Nambiar and Poess, 2006). The characteristics of the workloads used are summarized in Table 2. Training data includes pairs of queries and plans, encoded to be consumable by an ML model, as described in the following.
CEB 1000 queries are randomly sampled from various query templates in the CEB with up to 12 joins over 21 tables from the IMDB dataset. These are split to non-overlapping subsets of 800, 100, and 100 queries for training, validation, and test sets. Each experiment is repeated 5 times with different random seeds and the average results are reported.
@@
TPC-DS dataset is used as the third benchmark. The original benchmark includes only referential integrity (RI) joins, hence star-schema and snowflake-schema queries only. It does not include many-to-many joins or cyclic joins. In order to enforce more challenging scenarios, we extend this benchmark by adding a) many-to-many joins, b) cyclic joins, c) local predicate on new columns that are correlated with join predicate columns. This extended dataset is used to randomly generate 13k synthetic queries with various query templates and join patterns. The characteristics of these queries are controlled by parameters that a query generator takes as input, such as the number of joins, join types (inner-join, outer-join, left/right outer join, anti-join), number of join and local predicates, join and local predicate operator types (==, ¡, ¿, ¡=, ¿=), etc. Each query has up to 4 joins with each join having up to 3 join predicates and each table with up to 5 local predicates. The generated queries are uniformly distributed across different numbers of joins, such that 1-join queries are as frequent as 2-, 3-, and 4-join queries. The join predicates are uniformly sampled from the referential integrity joins. The queries may exhibit various join graphs (linear, star, snowflake, cyclic). The predicates used for each join as well as the local predicates defined on each base table may have correlations to simulate more realistic scenarios. The queries may also exhibit join-crossing correlations between the predicate columns across different joins and local predicates. As such, this workload ensures sufficient randomness and complexity in the query patterns to represent more challenging scenarios for generalization. 12k of the queries are used for training, 500 for validation, and 500 for testing. Each experiment is repeated five times with different random seeds and the average results are reported.
@@
Workload Queries Max Joins Dataset Tables
CEB 1000 12 IMDB 21
JOB 113 28 IMDB 21
TPC-DS 13000 4 Synthetic 24
Plan Generation Each query is compiled with various hint sets and the different plans generated by the optimizer are captured. The first compilation is done without any constraints, giving the default plan chosen by the optimizer. Additionally, each query is compiled using 13 hint sets (see Appendix A.3). Hint sets produce locally optimal plans that are expected to be better than the optimizer’s plan in some scenarios, as shown by Bao (Marcus et al., 2024). Although this does not represent the way the optimizer traverses the search space, it is sufficient for providing a diversified set of potentially optimal plans for the purpose of this study.
Collecting Labels Each query is executed using each of the guidelines and the execution time is measured. A timeout threshold is used to terminate too long executions. This threshold is dynamically set to be 3 times the time it takes to run the query using the optimal plan, rounded up to the nearest greater integer.
Preprocessing Preprocessing is performed for all splits based on the statistics obtained from the training data. The input data is preprocessed by performing null imputation and min-max scaling. Null values are replaced by the mean value of the respective features. Then min-max scaling brings the range of the values for each feature between zero and one: $X^{t}=X_{scaled}=\frac{{X-\min(X)}}{{\max(X)-\min(X)}}$. The labels are pre-processed by applying logarithmic transformation and min-max scaling. The execution times can range greatly and typically exhibit a large skewness towards zero. Log transformation is applied to the labels to reduce skewness. This is performed with the following formula, where $y_{log}$ represents the transformed label: $y_{log}=\log_{10}(y)$. The Min-Max scaling then brings the range of values between zero and one. This is a desirable range for deep learning models that use the Sigmoid activation function in their output layers: $y^{t}=y_{scaled}=\frac{{y_{log}-\min(y_{log})}}{{\max(y_{log})-\min(y_{log})}}$.
@@
To avoid over-fitting, three mechanisms are used in model training: early stopping, dropout, and reducing learning rate on plateau. Applying dropout at every layer is additionally necessary to enable the variational inference needed to obtain model uncertainty. The dropout rate along with other architectural parameters such as the number of layers, number of neurons, and learning rate are tuned using the Asynchronous Successive Halving Algorithm (Li et al., 2020). This algorithm allows for exploring a large number of parameter combinations while limiting the total amount of time needed for tuning.

### v3

@@ 3.2.2. Risk Quantification
Suboptimality Risk As stated in the problem formulation with Approach 1, the goal of robust query optimization is to find a plan with minimum risk of suboptimality. This can be formalized as the likelihood of the plan being suboptimal compared with the alternative plan(s). To explain this, let us revisit Example 3.2. We assume that the cost of the optimal plans in all scenarios is 8, and it is 10 for the alternative plans. The difference between the scenarios is the standard deviation of the cost distributions. Let us assume the cost for the optimal plan (plan X) is denoted by $C(X)\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and the cost for the alternative plan (plan Y) is denoted by $C(Y)\sim\mathcal{N}(\mu_{y},\sigma_{y}^{2})$. Also, let us assume the covariance between $C(X)$ and $C(Y)$ is denoted as $\text{Cov}(x,y)$. Then the probability of the alternative plan being cheaper than the optimal plan is denoted by:
@@
An accurate estimation of the risk associated with selecting the plan $X$ over the plan $Y$ would require the probability distribution of the cost for each plan, as well as the pairwise joint distribution $C(X)\cdot C(Y)$. Assuming independence between $C(X)$ and $C(Y)$ results in overestimating the risk by increasing the variance of $C(X)-C(Y)$. We assume independence for two reasons: a) the computation is impractical and the overhead is prohibitive, and b) even with the independence assumption made, the risk quantification provides significant improvements. The effectiveness of the simplified quantification, as demonstrated in the experimental study, can be explained by the fact that the overestimation effect for all pairs of plans is consistent, effectively canceling each other out.
The Probability Distribution The cost and execution time are not inherently normally distributed variables, as assumed by the loss function (Equation 4). However, the targets can be transformed so that they approximately conform to a normal distribution. This is handled in the preprocessing phase as explained in Section 5.1. The impact of the transformation on the target values is shown in Figure3, where the latency distributions of the JOB benchmark before and after the preprocessing step are shown.
The above discussions support that the assumptions made in Roq are both justified and inconsequential as stated in the design objective (b) in the introduction.
Figure 3. Target distributions before and after transformation
@@
3:  for $p_{i}\in\mathcal{P}$ do
4:     $\Delta_{ij}\sim N(\mu_{i}-\mu_{j},\sigma_{i}^{2}+\sigma_{j}^{2})$ for all $j\neq i$
5:     $R(p_{i})=\frac{1}{S-1}\sum_{j\neq i}P(\Delta_{ij}>0)$
6:  end for
7:  $P^{*}=\arg\min_{p_{i}\in\mathcal{P}}R(p_{i})$
8:  return $P^{*}$
@@
3:  $P^{*}=\arg\min_{p_{i}\in\mathcal{P}}(\mu_{p_{i}}+f_{s}\cdot\sigma_{p_{i}})$
4:  return $P^{*}$
3.3. Risk-aware Plan Evaluation
Classic optimizers select a plan from a set of plans solely based on expected cost. They assume that the estimated expected costs have zero uncertainties. In contrast, when the uncertainties, and, thus, risk, are quantified, a new family of plan evaluation and selection methods can be devised that account for them. We propose three risk-aware plan evaluation strategies. While the first two use risk measures directly for selecting plans, the third uses risk to prune the search space. We design algorithms that implement these strategies.
3.3.1. Plan Selection by Suboptimality Risk
This strategy follows the first approach to the problem formulation where the problem is defined as a risk minimization one. The suboptimality risk is computed for every plan in the search space, and the one with the minimum risk is selected, based on the risk quantification method suggested in Section 3.2. Given a set $\mathcal{P}=\{p_{i}\,|\,\forall i\in\{1,\ldots,S\}\}$ of $S$ plans, it computes the suboptimality risk $R(p_{i})$ for every $p_{i}$. The plan with a minimum $R(p_{i})$ is selected. The algorithm can use either plan or estimation risk, or a combination of the two.
3.3.2. Conservative Plan Selection
This strategy assumes that each plan has a higher execution cost than its estimated cost, and this difference is proportional to the standard deviation of its estimated cost. It eliminates risky plans (i.e., plans with high standard deviation) that happen to be low cost. Given a set $\mathcal{P}=\{p_{i}\,|\,\forall i\in\{1,\ldots,S\}\}$ of $S$ plans, where each $p_{i}$ has a cost distribution $C_{p_{i}}\sim N(\mu_{p_{i}},\sigma_{p_{i}}^{2})$, rather than picking the plan with minimum $\mu_{p_{i}}$, it picks the plan with minimum $\mu_{p_{i}}+f_{s}\cdot\sigma_{p_{i}}$, where $f_{s}$ is a parameter tuned on a validation set. The algorithm can be specialized to use either the plan or the estimation risk, or both. It follows the second approach to the problem formulation where a linear combination of the two objectives $\alpha\cdot\mu_{p_{i}}+(1-\alpha)\cdot\sigma_{p_{i}}$ is minimized, where $f_{s}=\frac{1-\alpha}{\alpha}$.
@@ 3.3.3. Search Space Pruning by Plan and Estimation Risks
This strategy prunes the search space to eliminate highly risky plans, based on the plan or the estimation risks. Algorithm 3 implements this strategy: given a set $\mathcal{P}=\{p_{i}\mid\forall i\in\{1,\dots,S\}\}$ of $S$ plans along with their plan and estimation risks ($R_{p}$ and $R_{e}$ respectively), and fractions of plans to be pruned with the highest plan and estimation risks ($f_{pr}$ and $f_{er}$ respectively), thresholds $\varphi_{pr}$ and $\varphi_{er}$ are computed for plan and estimation risks, respectively. Any plan with either plan or estimation risks above the thresholds is eliminated. Parameters $f_{pr}$ and $f_{er}$ are tuned using the samples in the validation set. Choosing the plan from the remaining ones relies on a plan selection strategy.
@@
1:  Inputs: $\mathcal{P}=\{p_{i}\mid\forall i\in\{1,\dots,S\}\}$: a set of plans to be evaluated $R_{e}=\{R_{e}(p_{i})\mid\forall i\in\{1,\dots,S\}\}$: a set of estimation risks $R_{p}=\{R_{p}(p_{i})\mid\forall i\in\{1,\dots,S\}\}$: a set of plan risks $f_{er}$: a fraction of plans to be pruned by estimation risk $f_{pr}$: a fraction of plans to be pruned by plan risk
2:  Output: $P^{\prime}\subset\mathcal{P}$: pruned search space
3:  $\varphi_{er}=\text{sorted}(R_{e})[\lfloor|\mathcal{P}|\times f_{er}\rfloor]$
4:  $\varphi_{pr}=\text{sorted}(R_{p})[\lfloor|\mathcal{P}|\times f_{pr}\rfloor]$
5:  $P^{\prime}=\{p_{i}\in\mathcal{P}\mid R_{e}(p_{i})\leq\varphi_{er}\text{ and }R_{p}(p_{i})\leq\varphi_{pr}\}$
6:  return $P^{\prime}$
Figure 4. a) Architecture of the risk-aware learned cost model, including representation learning and estimator components. This architecture allows quantification of model and data uncertainties, b) Architecture of the extended transformerConv GNN model block that enables receiving and processing graph level attributes in addition to node and edge attributes
[…]
@@ 4. Risk-aware Learned Cost Model
We propose a robust learned cost model, which infers the execution time, and the plan, estimation, and suboptimality risks. Figure 4-a illustrates the model’s architecture. It takes as input queries and plans and predicts the execution time along with expected variance. It includes two main modules: The first learns the representation of query and plan combinations; the second takes this representation and learns the query-plan associations and their impact on runtime.
@@
Query Encoding. We use the query’s join graph to encode tables, local predicates, join predicates, aggregations, etc, similarly to previous work (Yu et al., 2022). In this work, we encode table statistics, such as cardinality and selectivity from local predicates and correlations between predicate columns (captured by the optimizer statistics) as node attributes. We incorporate characteristics of joins, such as join type, join predicate operator, join column skewness and join selectivity, as edge attributes. The optimizer statistics are extended to capture the join characteristics, too. In addition to node and edge attributes, we capture high-level characteristics of the join graph, such as the number of tables and joins, and the join graph topology, as global attributes. This representation is agnostic to the plan (join orders and plan operators) that will be used to execute the query.
@@
Plan Encoding. We use a vectorized plan tree that preserves the structure of the plan (Marcus et al., 2019). Each node corresponds to a plan operator and captures the operator type, including different access, join, and aggregate operators, using one-hot encoding. In addition, it captures the table(s) in the subgraph originating at the node using an embedding learned by the Query Processor (see Figure 4-a).
@@
Representation Learning. The representation learning module produces query embeddings (Query Processor) and plan embeddings (Plan Processor). Although we partially borrow the architecture of the Plan Processor from prior work (Marcus et al., 2019), the architecture of the Query Processor and the output that feeds the Plan Processor are novel and support the robustness of the proposed model.
@@
Query Processor. We use Graph Neural Nets (GNNs) to process the encoded join graph and learn query and table representations. Each GNN layer serves as a message passing step in which each table receives information from its adjacent tables, each additional layer adding to the depth of the message passing in the graph. This allows learning a representation for each table such that it contains information about its adjacent tables and the whole join graph. Additionally, we use mean and max pooling to aggregate the information from all tables and produce a representation for the whole join graph. Although there are various GNN implementations, we use TransformerConv (Shi et al., 2021) in our architecture. The TransformerConv uses an attention mechanism (Vaswani et al., 2017) to condition the amount of information each node receives from its neighboring nodes by the attributes of the target node, the neighboring nodes, and the edges connecting them to the target node. This is a suitable option for processing the join graphs as the amount of data processed through a subgraph should be conditioned by the tables and the joins involved. We extend TransformerConv as illustrated in Figure 4-b, to receive and process graph-level attributes, broadcast them to every node, and apply mean and max pooling to aggregate them back into a graph level embedding. This novel architecture allows us to effectively produce query embeddings.
@@
Plan Processor. The Plan Processor takes the vectorized plan tree as input and augments nodes with the embedding of the corresponding table(s) produced by the Query Processor. If a node involves multiple tables, their corresponding embeddings are aggregated using mean-pooling before being concatenated to the plan node. The augmented plan tree is then processed through multiple layers of Tree Convolutional Neural Networks (TCNN), ultimately aggregated into a one-dimensional vector using dynamic pooling (Mou et al., 2016). This architecture is proven to be effective in processing query plan trees by prior works (Marcus et al., 2019, 2024; Yu et al., 2022). In our tests, TCNN was superior to alternative architectures for processing trees such as GNNs or Tree-LSTMs (Yu et al., 2020). The embeddings generated by the final layer of this model serve as the representation of the query plan.
@@
Novelty of Roq’s Model Architecture. While some components of Roq’s model are derived from prior works, to the best of our knowledge, their composition and extension in this architecture towards learning cost and quantifying risks are novel. Specifically, the novelty lies in the adoption and extension of the state-of-the-art GNNs to learn generalizable representations for tables and join graphs and their integration into TCNNs, as well as the employment of two separate branches of MLPs for predicting the plan cost and risk simultaneously.
@@ 5. Experimental Study
We experimentally compare Roq with the state-of-the-art.
@@ 5.1. Experimental Setup
We use the following benchmarks: The Join Order Benchmark (JOB) (Leis et al., 2015), the Cardinality Estimation Benchmark (CEB) (Negi et al., 2021), and the Decision Support Benchmark (DSB) (Nambiar and Poess, 2006). The characteristics of the workloads used are summarized in Table 2. Training data includes pairs of queries and plans, encoded to be consumable by an ML model, as described in the following.
CEB 1000 queries are randomly sampled from query templates in CEB with up to 12 joins over 21 tables from the IMDB dataset. These are split to non-overlapping subsets of 800, 100, and 100 queries for training, validation, and test. Each experiment is repeated 5 times with different random seeds and the average results are reported.
@@
DSB 1000 queries are generated based on the Decision Support Benchmark (DSB) (Ding et al., 2021), an extension of TPC-DS that enhances data generation by introducing more realistic, skewed, and correlated data distributions rather than the mostly independent, uniform distributions in TPC-DS. It augments query templates with predicate filters for fine-grained data slicing and adds new templates featuring more complex join patterns (such as non-equi and many-to-many joins), resulting in many more distinct query instances. The queries are generated based on 15 distinct query templates and are split into 800, 100, and 100 queries for training, validation, and test sets, respectively. Each experiment is repeated 5 times with different random seeds and the average results are reported.
Plan Generation Each query is compiled with various hint sets and the different plans generated by the optimizer are captured. The first compilation is done without any constraints, leading to the use of the default plan chosen by the optimizer. Additionally, each query is compiled using 13 hint sets (see Appendix A.3). Hint sets produce locally optimal plans that are expected to be better than the optimizer’s plan in some scenarios, as shown by Bao (Marcus et al., 2024). Although this does not represent the way the optimizer traverses the search space, it is sufficient for providing a diversified set of potentially optimal plans for the purpose of this study.
@@
Queries Templates Max Joins Dataset Tables
CEB 1000 16 12 IMDB 21
JOB 113 33 28 IMDB 21
DSB 1000 15 13 TPC-DS 24
Collecting Labels Each query is executed using each of the guidelines and the execution time is measured. A timeout threshold is used to terminate too long executions. This threshold is dynamically set to be 10 times the time it takes to run the query using the best performing plan found so far, rounded up to the nearest greater integer. This threshold allows a higher level of exploration to obtain information on more risky plans.
Preprocessing Preprocessing is performed for all splits based on the statistics obtained from the training data. Min-max scaling is applied to bring the range of the values for each feature between zero and one: $X_{t}=\frac{{X-\min(X)}}{{\max(X)-\min(X)}}$. We apply logarithmic transformation and min-max scaling to labels. The execution times can range greatly and typically exhibit a large skewness towards zero. Log transformation is applied to the labels to reduce skewness. This is performed with the following formula, where $y_{log}$ represents the transformed label: $y_{log}=\log_{10}(y)$. The Min-Max scaling then brings the range of values between zero and one. This is a desirable range for deep learning models that use the Sigmoid activation function in their output layers: $y_{t}=\frac{{y_{log}-\min(y_{log})}}{{\max(y_{log})-\min(y_{log})}}$.
@@
To avoid over-fitting, three mechanisms are used in training: early stopping, dropout, and reducing learning rate on plateau. Applying dropout at every layer is necessary to enable the variational inference needed to obtain model uncertainty. The dropout rate along with other parameters such as the number of layers, number of neurons, and learning rate are tuned using the Asynchronous Successive Halving Algorithm (Li et al., 2020). This allows for exploring a large number of parameter combinations while limiting the total time needed for tuning.